# Multi-SpaCE: Multi-Objective Subsequence-based Sparse Counterfactual Explanations for Multivariate Time Series Classification

### v1

@@ 5 Conclusions and Future Work
Although this work addresses an important gap, there are several promising directions for future improvements. One area is the use of generative models, such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs) or Diffusion Models, to create richer and more realistic modifications of the original instance, rather than relying on substitutions from the NUN. This would improve plausibility and allow for greater diversity in counterfactuals, particularly in complex datasets. Another promising direction is the inclusion of amortized generation techniques, which could leverage knowledge from previously explained instances to generate counterfactual explanations. This would reduce the computational overhead associated with optimization-based methods, enabling faster and more efficient generation.
@@
Appendix A Model details
A.1 Black-box Classifier
We trained three different architectures for each dataset to better adapt the models to the specific characteristics of the data, serving as a lighter alternative to grid search or Bayesian optimization. Each model randomly splits 10% of the training data for validation and is trained over 200 epochs, using early stopping with a patience of 30 epochs. Additionally, a learning rate reduction is applied when the validation loss plateaus, with a patience of 10 epochs. The initial learning rate is set to 0.001, the batch size to 32, and a dropout rate of 20% is applied across all models. All the architectures are based on convolutional layers. The architectures differ based on the following parameters:
  • CNN-simple: Two convolutional layers with 32 and 64 channels, each with a kernel size of 5, followed by a Dense layer with 64 units.
  • CNN-intermediate: Three convolutional layers with 16, 32, and 64 channels, with kernel sizes of 5, 5, and 3 respectively, followed by a Dense layer with 64 units.
  • CNN-complex: Four convolutional layers with 16, 32, 64, and 128 channels, with kernel sizes of 5, 5, 5, and 3 respectively, followed by a Dense layer with 128 units.
Tables 6 and LABEL:tab:classifiers_uni present the results for each architecture on the multivariate and univariate datasets, respectively. For each dataset, the best-performing model is highlighted in bold and is selected as the classifier to be explained. The horizontal line in the middle of each table represents the 80% F1-score threshold used to ensure explanations are generated only for datasets with sufficiently accurate classifiers.
Dataset CNN-simple CNN-intermediate CNN-complex
PenDigits 0.98 0.98 0.98
SelfRegulationSCP1 0.87 0.9 0.85
NATOPS 0.89 0.87 0.84
Epilepsy 0.76 0.88 0.24
PEMS-SF 0.86 0.87 0.8
ArticularyWordRecognition 0.86 0.85 0.85
UWaveGestureLibrary 0.85 0.81 0.82
Cricket 0.84 0.8 0.51
RacketSports 0.84 0.71 0.69
BasicMotions 0.84 0.71 0.67
FaceDetection 0.66 0.65 0.64
Heartbeat 0.64 0.63 0.65
FingerMovements 0.55 0.55 0.55
HandMovementDirection 0.53 0.49 0.47
SelfRegulationSCP2 0.45 0.49 0.5
MotorImagery 0.47 0.48 0.49
LSST 0.3 0.3 0.37
StandWalkJump 0.31 0.36 0.17
EthanolConcentration 0.25 0.13 0.16
Handwriting 0.2 0.18 0.17
Libras 0.01 0.01 0.03
Table 6: Black-box classifier F1-scores on multivariate test sets.
Table 7: Black-box classifier F1-scores on univariate test sets.
Dataset CNN-simple CNN-intermediate CNN-complex
TwoPatterns 0.99 1.0 1.0
Wafer 0.99 0.99 0.99
SyntheticControl 0.98 0.98 0.97
Strawberry 0.97 0.98 0.98
Plane 0.33 0.05 0.98
StarLightCurves 0.96 0.96 0.97
ItalyPowerDemand 0.96 0.96 0.33
CinCECGTorso 0.95 0.77 0.71
GunPoint 0.33 0.41 0.95
Coffee 0.29 0.44 0.93
HandOutlines 0.92 0.92 0.89
FordA 0.9 0.89 0.92
NonInvasiveFatalECGThorax2 0.9 0.91 0.91
SwedishLeaf 0.87 0.88 0.91
ECGFiveDays 0.91 0.62 0.88
Trace 0.25 0.9 0.06
ECG5000 0.9 0.9 0.9
ProximalPhalanxOutlineCorrect 0.9 0.89 0.89
FaceFour 0.9 0.89 0.78
CBF 0.9 0.89 0.34
NonInvasiveFatalECGThorax1 0.89 0.89 0.9
Fish 0.21 0.88 0.23
MoteStrain 0.87 0.85 0.86
ProximalPhalanxOutlineAgeGroup 0.81 0.85 0.86
ECG200 0.85 0.31 0.23
FacesUCR 0.84 0.85 0.84
FaceAll 0.84 0.71 0.72
SonyAIBORobotSurface2 0.83 0.81 0.79
MiddlePhalanxOutlineCorrect 0.77 0.81 0.8
PhalangesOutlinesCorrect 0.81 0.81 0.81
ChlorineConcentration 0.8 0.8 0.79
ToeSegmentation2 0.76 0.8 0.09
DistalPhalanxOutlineCorrect 0.79 0.71 0.73
Yoga 0.37 0.63 0.79
BeetleFly 0.44 0.79 0.44
FordB 0.7 0.7 0.75
ProximalPhalanxTW 0.74 0.74 0.73
DistalPhalanxOutlineAgeGroup 0.74 0.63 0.73
Lightning2 0.43 0.69 0.73
MedicalImages 0.69 0.69 0.71
Ham 0.71 0.68 0.71
Earthquakes 0.66 0.7 0.68
SonyAIBORobotSurface1 0.68 0.61 0.43
Lightning7 0.6 0.66 0.68
TwoLeadECG 0.33 0.33 0.66
FiftyWords 0.58 0.6 0.65
DistalPhalanxTW 0.6 0.6 0.65
ElectricDevices 0.65 0.65 0.64
ShapesAll 0.54 0.6 0.63
InsectWingbeatSound 0.62 0.6 0.62
Symbols 0.62 0.6 0.5
SmallKitchenAppliances 0.55 0.59 0.6
Computers 0.57 0.6 0.6
Adiac 0.59 0.04 0.06
BirdChicken 0.58 0.52 0.33
WordSynonyms 0.53 0.52 0.57
LargeKitchenAppliances 0.52 0.56 0.55
DiatomSizeReduction 0.56 0.46 0.13
ToeSegmentation1 0.56 0.46 0.53
MiddlePhalanxTW 0.54 0.52 0.33
MiddlePhalanxOutlineAgeGroup 0.53 0.46 0.52
WormsTwoClass 0.49 0.51 0.42
ShapeletSim 0.5 0.45 0.35
RefrigerationDevices 0.45 0.48 0.48
Herring 0.44 0.44 0.44
Worms 0.37 0.28 0.4
ScreenType 0.37 0.38 0.33
Wine 0.33 0.33 0.33
Beef 0.33 0.07 0.07
Haptics 0.33 0.2 0.14
OSULeaf 0.29 0.22 0.15
Car 0.17 0.09 0.26
Mallat 0.1 0.24 0.05
OliveOil 0.23 0.03 0.14
Meat 0.17 0.17 0.17
InlineSkate 0.15 0.1 0.07
ArrowHead 0.14 0.14 0.14
Phoneme 0.07 0.08 0.05
A.2 Autoencoder
@@
Tables 8 and 9 present the results for each architecture on the multivariate and univariate datasets, respectively.
@@ A.3 Isolation Forest
The model with the highest silhouette score on the test set is selected. The chosen model is then used to calculate the Outlier Score, reported in Table 4 and Table 5. This process is performed independently for each dataset.
A.4 Local Outlier Factor
@@
The model with the highest silhouette score on the test set is selected. The chosen model is then used to calculate the Outlier Score, reported in Table 4 and Table 5. This process is performed independently for each dataset.
Appendix B Mask and mutation ablation study
B.1 Extension/Compression mutation probability
The first experiment investigated the impact of extension and compression mutation probabilities under the common and independent mask settings. The objective was to identify optimal probabilities and assess the sensitivity of the optimization process to this parameter. The tested probabilities were $p^{e}=p^{c}=[0.1,0.25,0.5,0.75,0.9]$, while $p^{p}$ was fixed at 0. The NSGA-II parameters were set to a population size of $N=100$ and a total of $G=100$ generations.
In terms of validity, all configurations achieved a perfect score. Regarding proximity, Table 10 shows that extreme probability values (both low and high) perform best for the common mask setting, whereas higher values are preferable for the independent mask setting. Regarding plausibility, as shown in Table 11, intermediate probabilities yield better results for both mask types.
Dataset Common mask
$p^{e}=p^{c}=0.1$ $p^{e}=p^{c}=0.25$ $p^{e}=p^{c}=0.5$ $p^{e}=p^{c}=0.75$ $p^{e}=p^{c}=0.9$
AWR 16.02 16.45 16.67 16.4 15.93
BasicMotions 61.28 60.81 63.1 61.92 60.59
Cricket 51.87 51.26 51.04 51.12 50.1
Epilepsy 9.6 9.98 10.09 9.73 9.21
NATOPS 9.37 9.45 9.43 9.46 9.42
PenDigits 45.75 45.89 46.06 45.97 46.31
RacketSports 51.94 52.33 52.55 53.2 53.98
SR-SCP1 401.97 449.21 462.9 446.12 409.22
UWave 15.7 16.44 16.75 16.26 15.4
Average Rank 2.0 3.33 4.11 3.44 2.11
Dataset Independent mask
$p^{e}=p^{c}=0.1$ $p^{e}=p^{c}=0.25$ $p^{e}=p^{c}=0.5$ $p^{e}=p^{c}=0.75$ $p^{e}=p^{c}=0.9$
AWR 16.71 16.8 16.47 16.53 16.93
BasicMotions 66.23 64.77 65.79 65.94 65.64
Cricket 53.82 54.16 52.79 52.66 53.77
Epilepsy 9.66 10.05 10.09 9.63 9.51
NATOPS 8.4 8.11 8.06 7.8 7.63
PenDigits 45.06 45.25 45.33 45.44 45.8
RacketSports 41.69 41.22 41.09 40.36 40.34
SR-SCP1 545.23 527.98 515.13 520.43 507.87
UWave 14.85 15.07 14.99 14.96 14.43
Average Rank 3.67 3.67 2.89 2.56 2.22
Dataset Common mask
$p^{e}=p^{c}=0.1$ $p^{e}=p^{c}=0.25$ $p^{e}=p^{c}=0.5$ $p^{e}=p^{c}=0.75$ $p^{e}=p^{c}=0.9$
AWR 0.69 0.68 0.68 0.68 0.69
BasicMotions 0.29 0.29 0.28 0.29 0.3
Cricket 0.78 0.73 0.73 0.73 0.79
Epilepsy 0.6 0.59 0.59 0.59 0.61
NATOPS 0.57 0.57 0.57 0.57 0.57
PenDigits 0.28 0.28 0.28 0.28 0.27
RacketSports 0.68 0.68 0.67 0.67 0.67
SR-SCP1 0.45 0.44 0.44 0.44 0.44
UWave 0.66 0.61 0.61 0.63 0.67
Average Rank 3.33 1.56 1.11 1.44 3.11
Dataset Independent mask
$p^{e}=p^{c}=0.1$ $p^{e}=p^{c}=0.25$ $p^{e}=p^{c}=0.5$ $p^{e}=p^{c}=0.75$ $p^{e}=p^{c}=0.9$
AWR 0.8 0.76 0.75 0.76 0.81
BasicMotions 0.33 0.34 0.33 0.33 0.33
Cricket 0.96 0.85 0.82 0.93 1.06
Epilepsy 0.64 0.62 0.61 0.62 0.63
NATOPS 0.74 0.72 0.7 0.72 0.74
PenDigits 0.27 0.27 0.27 0.27 0.27
RacketSports 0.74 0.74 0.74 0.74 0.74
SR-SCP1 0.53 0.52 0.5 0.51 0.52
UWave 0.8 0.68 0.62 0.66 0.93
Average Rank 3.22 2.33 1.0 1.78 3.22
This trend is also reflected in the Sparsity-NoS average metric, where intermediate probabilities provide the best balance for both mask settings, as depicted in Figure 3.
(a) Common mask.
(b) Independent Mask.
B.2 Pruning mutation probability
In this experiment, we introduced pruning mutation into the configurations identified in the previous section. Specifically, we tested pruning mutation probabilities $p^{p}=[0.05,0.1,0.2,0.35,0.5]$ and also evaluated the case without pruning mutation. The NSGA-II parameters remained unchanged ($N=100$ and $G=100$).
In terms of validity, all configurations achieved perfect scores. For proximity, as shown in Table 12, lower pruning probabilities generally performed better for both mask settings, with the absence of pruning mutation leading to slightly worse proximity results. However, looking at the absolute differences, they were not significant. Similar trends were observed for plausibility in Table 13, where the absence of pruning mutation occasionally improved results in the independent mask setting.
Dataset Common mask ($p^{e}=p^{c}=0.75$)
$p^{p}=0$ $p^{p}=0.05$ $p^{p}=0.1$ $p^{p}=0.2$ $p^{p}=0.35$ $p^{p}=0.5$
AWR 16.4 16.31 16.31 16.3 16.28 16.92
BasicMotions 61.92 61.85 63.93 62.57 63.5 63.28
Cricket 51.12 48.46 50.35 51.29 52.88 55.08
Epilepsy 9.73 9.7 9.74 9.66 9.66 9.9
NATOPS 9.46 9.42 9.41 9.39 9.44 9.35
PenDigits 45.97 45.91 46.3 45.96 46.17 46.12
RacketSports 53.2 52.6 52.58 52.35 52.91 53.53
SR-SCP1 446.12 442.06 437.94 467.2 514.05 555.71
UWave 16.26 15.34 15.45 15.27 14.98 15.47
Average Rank 4.11 2.33 3.56 2.33 3.56 4.89
Dataset Independent mask
$p^{p}=0$ $p^{p}=0.05$ $p^{p}=0.1$ $p^{p}=0.2$ $p^{p}=0.35$ $p^{p}=0.5$
AWR 16.47 16.4 16.2 16.61 17.37 17.89
BasicMotions 65.79 66.07 66.71 66.73 66.78 68.11
Cricket 52.79 50.66 51.21 52.06 53.78 56.12
Epilepsy 10.09 9.92 9.96 9.78 10.18 10.67
NATOPS 8.06 7.41 7.35 7.12 7.6 7.71
PenDigits 45.33 45.06 45.32 45.17 45.02 44.88
RacketSports 41.09 40.62 40.92 39.58 40.45 40.93
SR-SCP1 515.13 495.68 504.29 512.49 561.29 589.42
UWave 14.99 14.74 14.61 14.56 14.68 15.53
Average Rank 4.33 2.33 2.67 2.44 4.0 5.22
Table 12: Proximity results for Common/Independent mask settings using different pruning mutation probabilities.
Dataset Common mask ($p^{e}=p^{c}=0.75$)
$p^{p}=0$ $p^{p}=0.05$ $p^{p}=0.1$ $p^{p}=0.2$ $p^{p}=0.35$ $p^{p}=0.5$
AWR 0.68 0.68 0.68 0.68 0.68 0.73
BasicMotions 0.29 0.29 0.27 0.28 0.28 0.28
Cricket 0.73 0.72 0.84 0.98 1.08 1.12
Epilepsy 0.59 0.59 0.59 0.59 0.62 0.66
NATOPS 0.57 0.57 0.57 0.57 0.57 0.57
PenDigits 0.28 0.28 0.27 0.28 0.27 0.28
RacketSports 0.67 0.67 0.67 0.68 0.67 0.67
SR-SCP1 0.44 0.44 0.45 0.48 0.51 0.53
UWave 0.63 0.61 0.6 0.64 0.85 1.34
Average Rank 2.0 1.78 1.44 2.89 2.89 4.11
Dataset Independent mask ($p^{e}=p^{c}=0.5$)
$p^{p}=0$ $p^{p}=0.05$ $p^{p}=0.1$ $p^{p}=0.2$ $p^{p}=0.35$ $p^{p}=0.5$
AWR 0.75 0.77 0.78 0.82 0.83 0.85
BasicMotions 0.33 0.35 0.35 0.36 0.35 0.34
Cricket 0.82 0.95 1.03 1.07 1.09 1.13
Epilepsy 0.61 0.62 0.62 0.66 0.68 0.71
NATOPS 0.7 0.72 0.73 0.76 0.78 0.79
PenDigits 0.27 0.27 0.27 0.26 0.27 0.27
RacketSports 0.74 0.74 0.74 0.74 0.74 0.74
SR-SCP1 0.5 0.53 0.53 0.54 0.53 0.54
UWave 0.62 0.63 0.74 0.93 1.54 1.93
Average Rank 1.11 2.0 2.44 3.67 3.67 4.44
Table 13: Plausibility results for Common/Independent mask settings using different pruning mutation probabilities.
For the Sparsity-NoS average metric, Figure 4 shows that not using pruning mutation generally yields better results for both mask settings. Considering these findings, and the additional computational cost introduced by pruning mutation, we decided to exclude pruning mutation during the same stage as extension/compression mutation. High pruning probabilities often led to non-essential changes and complicated convergence.
(a) Common mask.
(b) Independent Mask.
Figure 4: Arithmetic mean between normalized sparsity and normalized number of subsequences using different pruning mutation probabilities.
B.3 Final iterations using only pruning mutation
Although pruning mutation was excluded during earlier stages of optimization, we hypothesized that applying it during a final optimization stage could improve contiguity and eliminate unnecessary subsequences. To test this, we divided the optimization process into two stages, as implemented in the final version of Multi-SpaCE described in Section 3.
The first stage consisted of $G_{1}=75$ generations with the best parameters identified earlier: $p^{e}=p^{c}=0.75$ and $p^{p}=0$ for the common mask setting, and $p^{e}=p^{c}=0.5$ and $p^{p}=0$ for the independent mask setting. In the second stage, the resulting Pareto front was further optimized over 25 generations ($G_{2}=25$) using $p^{e}=p^{c}=0$ and pruning probabilities $p^{p}=[0.25,0.5,0.75]$. During this stage, the independent mask setting was always used, regardless of the mask setting in the first stage.
In terms of validity, all configurations achieved perfect scores. For proximity, plausibility, and the Sparsity-NoS average, the best results were consistently achieved with the common mask setting for the first 75 generations, followed by independent mask pruning with a probability of $p^{p}=0.5$ during the final 25 iterations. This is the version used for the experiments in Section 4.
Dataset Common ($p^{e}=p^{c}=0.75$) Independent ($p^{e}=p^{c}=0.5$)
$p^{p}=0.25$ $p^{p}=0.5$ $p^{p}=0.75$ $p^{p}=0.25$ $p^{p}=0.5$ $p^{p}=0.75$
AWR 15.45 15.14 15.76 15.91 15.99 16.27
BasicMotions 63.86 62.91 61.98 64.7 64.17 65.05
Cricket 47.34 47.36 49.83 48.86 50.39 52.15
Epilepsy 9.63 9.44 9.39 9.79 9.8 9.93
NATOPS 8.75 7.95 7.72 7.47 7.07 7.38
PenDigits 45.35 45.11 45.47 45.4 45.69 45.61
RacketSports 42.15 41.65 41.25 40.05 40.43 40.01
SR-SCP1 429.04 428.47 438.52 479.62 483.24 495.55
UWave 14.4 14.42 14.59 14.63 14.59 14.47
Average Rank 2.89 2.33 3.11 3.78 4.22 4.56
Dataset Common ($p^{e}=p^{c}=0.75$) Independent ($p^{e}=p^{c}=0.5$)
$p^{p}=0.25$ $p^{p}=0.5$ $p^{p}=0.75$ $p^{p}=0.25$ $p^{p}=0.5$ $p^{p}=0.75$
AWR 0.69 0.7 0.69 0.75 0.75 0.75
BasicMotions 0.33 0.34 0.32 0.36 0.36 0.35
Cricket 0.76 0.76 0.75 0.84 0.84 0.85
Epilepsy 0.61 0.61 0.61 0.61 0.61 0.61
NATOPS 0.62 0.65 0.63 0.72 0.71 0.7
PenDigits 0.27 0.27 0.27 0.27 0.27 0.27
RacketSports 0.73 0.73 0.73 0.74 0.74 0.74
SR-SCP1 0.45 0.44 0.44 0.52 0.52 0.52
UWave 0.61 0.62 0.63 0.62 0.62 0.63
Average Rank 1.44 1.89 1.56 3.44 3.33 3.67

### v2

@@ 5 Conclusions and Future Work
Although this work addresses an important gap in time series explainability using counterfactuals, there are several promising directions for future improvements. One area is the use of generative models, such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs) or Diffusion Models, to create richer and more realistic modifications of the original instance, rather than relying on substitutions from the NUN. This would improve plausibility and allow for greater diversity in the counterfactuals, particularly in complex datasets. Another promising direction is the inclusion of amortized generation techniques, which could leverage knowledge from previously explained instances to generate counterfactual explanations. This would reduce the computational overhead associated with optimization-based methods, enabling faster and more efficient generation.
@@
Appendix A Mask and mutation ablation study
A.1 Extension/Compression mutation probability
The first experiment investigated the impact of extension and compression mutation probabilities under the common and independent mask settings. The objective was to identify optimal probabilities and assess the sensitivity of the optimization process to this parameter. The tested probabilities were $p^{e}=p^{c}\in\{0.1,0.25,0.5,0.75,0.9\}$, while $p^{p}$ was fixed at 0. The NSGA-II parameters were set to a population size of $N=100$ and a total of $G=100$ generations.
In terms of validity, all configurations achieved a perfect score. Regarding proximity, Table 6 shows that higher values perform best for the common mask setting, whereas higher intermediate values are preferable for the independent mask setting. Regarding plausibility, as shown in Table 7, intermediate probabilities yield better results for both mask types.
Dataset Common mask
$p^{e}=p^{c}=0.1$ $p^{e}=p^{c}=0.25$ $p^{e}=p^{c}=0.5$ $p^{e}=p^{c}=0.75$ $p^{e}=p^{c}=0.9$
AWR 19.3 19.38 19.24 19.26 19.2
BasicMotions 68.39 70.1 68.66 68.23 68.02
Cricket 57.38 54.68 53.5 55.53 58.99
Epilepsy 14.06 14.13 14.22 14.03 13.93
NATOPS 9.57 9.56 9.56 9.54 9.52
RacketSports 61.55 61.69 61.62 61.71 61.65
SR-SCP1 416.63 422.63 417.54 410.61 424.47
UWave 17.43 17.7 17.66 17.41 16.78
Average Rank 3.12 4.0 3.0 2.5 2.25
Dataset Independent mask
$p^{e}=p^{c}=0.1$ $p^{e}=p^{c}=0.25$ $p^{e}=p^{c}=0.5$ $p^{e}=p^{c}=0.75$ $p^{e}=p^{c}=0.9$
AWR 19.18 17.92 17.86 18.03 18.65
BasicMotions 71.71 70.5 68.4 69.9 70.28
Cricket 64.93 62.15 61.99 63.04 66.47
Epilepsy 13.49 13.37 13.51 13.46 13.62
NATOPS 8.97 8.86 8.52 8.46 8.59
RacketSports 51.23 50.15 49.9 49.56 48.87
SR-SCP1 465.32 419.69 394.73 386.52 406.45
UWave 16.17 16.23 15.68 15.95 16.15
Average Rank 4.5 3.25 1.88 2.0 3.38
Dataset Common mask
$p^{e}=p^{c}=0.1$ $p^{e}=p^{c}=0.25$ $p^{e}=p^{c}=0.5$ $p^{e}=p^{c}=0.75$ $p^{e}=p^{c}=0.9$
AWR 0.65 0.64 0.64 0.65 0.65
BasicMotions 0.27 0.26 0.26 0.26 0.27
Cricket 0.73 0.69 0.67 0.68 0.8
Epilepsy 0.56 0.54 0.53 0.54 0.56
NATOPS 0.57 0.57 0.57 0.57 0.57
RacketSports 0.62 0.62 0.62 0.61 0.62
SR-SCP1 0.43 0.43 0.43 0.43 0.44
UWave 0.62 0.6 0.59 0.6 0.63
Average Rank 2.88 1.62 1.12 1.62 3.62
Dataset Independent mask
$p^{e}=p^{c}=0.1$ $p^{e}=p^{c}=0.25$ $p^{e}=p^{c}=0.5$ $p^{e}=p^{c}=0.75$ $p^{e}=p^{c}=0.9$
AWR 0.75 0.71 0.7 0.71 0.77
BasicMotions 0.32 0.33 0.33 0.33 0.33
Cricket 0.94 0.85 0.84 0.93 1.05
Epilepsy 0.59 0.57 0.56 0.57 0.59
NATOPS 0.71 0.68 0.69 0.67 0.69
RacketSports 0.7 0.71 0.71 0.71 0.72
SR-SCP1 0.52 0.51 0.51 0.52 0.53
UWave 0.77 0.65 0.63 0.63 0.77
Average Rank 3.25 2.0 1.5 2.0 4.12
This behavior is also reflected in the Sparsity-NoS average metric, where intermediate probabilities provide the best balance for both mask settings, as depicted in Figure 5.
(a) Common mask.
(b) Independent Mask.
A.2 Pruning mutation probability
In this experiment, we introduced a pruning mutation into the configurations identified in the previous section, $p^{e}=p^{c}=0.75$ for the common mask, and $p^{e}=p^{c}=0.5$ for the independent mask. Specifically, we tested the pruning mutation probabilities $p^{p}\in\{0.05,0.1,0.2,0.35,0.5\}$ and also evaluated the case without pruning mutation. The conclusion was clear: not using the pruning mutation yields significantly better results for both mask settings. Considering these findings and the additional computational cost introduced by pruning mutation, we decided to exclude pruning mutation during the same stage as extension/compression mutation.
A.3 Final iterations using only pruning mutation
Although the pruning mutation was excluded during the earlier stages of the optimization, we hypothesized that applying it during a final optimization stage could improve contiguity and eliminate unnecessary subsequences. To test this, we divided the optimization process into two stages, as implemented in the final version of Multi-SpaCE shown in Algorithm 1.
The first stage consisted of $G_{1}=75$ generations with the best parameters identified earlier: $p^{e}=p^{c}=0.75$ and $p^{p}=0$ for the common mask setting, and $p^{e}=p^{c}=0.5$ and $p^{p}=0$ for the independent mask setting. In the second stage, the resulting Pareto front was further optimized over 25 generations ($G_{2}=25$) using $p^{e}=p^{c}=0$ and pruning probabilities $p^{p}\in\{0.25,0.5,0.75\}$. During this stage, the independent mask setting was always used, regardless of the mask setting in the first stage.
In terms of validity, all configurations achieved perfect scores. Regarding plausibility, the best results were consistently achieved with the common mask with final pruning $p^{p}=0.75$. For proximity, the common mask with final pruning $p^{p}=0.5$ performed slightly better than the $p^{p}=0.75$ setting. For the Sparsity-NoS average, common mask with both final pruning $p^{p}=0.5$ and $p^{p}=0.75$ achieved comparable results. Because of the difference in terms plausibility, $p^{p}=0.75$ is the version used for the experiments in Section 4.
Dataset Common ($p^{e}=p^{c}=0.75$) Independent ($p^{e}=p^{c}=0.5$)
$p^{p}=0.25$ $p^{p}=0.5$ $p^{p}=0.75$ $p^{p}=0.25$ $p^{p}=0.5$ $p^{p}=0.75$
AWR 17.16 16.33 16.52 17.09 16.51 17.12
BasicMotions 67.95 67.34 66.71 68.71 68.27 68.3
Cricket 48.64 48.85 51.67 54.6 59.62 60.81
Epilepsy 13.01 12.96 13.04 13.2 13.08 13.21
NATOPS 8.7 8.46 8.38 8.2 7.98 8.12
RacketSports 50.47 50.13 50.97 48.84 48.9 48.69
SR-SCP1 330.68 288.91 285.74 362.82 365.2 368.46
UWave 15.19 15.29 15.2 15.57 15.36 15.1
Average Rank 3.5 2.62 3.0 4.25 3.62 4.0
Table 8: Proximity results for Common/Independent mask structures with different pruning mutation probabilities during the final 25 iterations.
Dataset Common ($p^{e}=p^{c}=0.75$) Independent ($p^{e}=p^{c}=0.5$)
$p^{p}=0.25$ $p^{p}=0.5$ $p^{p}=0.75$ $p^{p}=0.25$ $p^{p}=0.5$ $p^{p}=0.75$
AWR 0.66 0.66 0.66 0.7 0.7 0.71
BasicMotions 0.34 0.34 0.33 0.35 0.35 0.35
Cricket 0.72 0.73 0.71 0.85 0.88 0.88
Epilepsy 0.58 0.58 0.56 0.57 0.56 0.57
NATOPS 0.61 0.62 0.61 0.71 0.71 0.71
RacketSports 0.7 0.7 0.69 0.71 0.71 0.71
SR-SCP1 0.46 0.46 0.45 0.51 0.51 0.51
UWave 0.59 0.59 0.61 0.61 0.62 0.63
Average Rank 2.0 2.38 1.25 3.75 3.88 4.5
Appendix B Model details
B.1 Black-box Classifier
We trained InceptionTime with default hyperparameters following the implementation in https://github.com/hfawaz/InceptionTime
B.2 Autoencoder
@@
Tables 10 and 11 present the results for each architecture on the multivariate and univariate datasets, respectively.
@@ B.3 Isolation Forest
The model with the highest silhouette score on the test set is selected. The chosen model is then used to calculate the Outlier Score on the main paper. This process is performed independently for each dataset.
B.4 Local Outlier Factor
@@
The model with the highest silhouette score on the test set is selected. The chosen model is then used to calculate the Outlier Score, on the main paper. This process is performed independently for each dataset.